Mark draws one card from a standard deck of 52. He wins $ 0.85 for the king of clubs, $ 0.40 for any other club, and $ 0.55 for any other a king. Compute Mark's expected winnings. Round your answer to the nearest thousandth.
Answer #1

let X be Mark winning

P(X=0.85) =P(king of club) =1/52

P(X=0.40) =P(other club) =12/52

P(X=0.55) =P(other king) =3/52

expected winnning =E(X) =xP(x) =0.85*(1/52)+0.40*(12/52)+0.55*(3/52)=0.140